pooja195 (pooja195):

Solve using the addition method 4x + 3y = -1 5x + 4y = 1 @hartnn

hartnn (hartnn):

since we need to use addition method, add those 2 equation and tell me what u get?

pooja195 (pooja195):

you cant add them you need to multiply and thats the part i dont get =/ I dont know what numbers to multiply

OpenStudy (anonymous):

no. add them first. i'll tell u wht to do after u done it

pooja195 (pooja195):

huh???

pooja195 (pooja195):

You cant you need opposites

OpenStudy (anonymous):

just do it...... add both equations....

OpenStudy (anonymous):

and u get something =0 right? then that and that must be equal....

pooja195 (pooja195):

9x+7y=0

OpenStudy (anonymous):

blah. it's so confusing when i say it that way

OpenStudy (anonymous):

light. so what do you know when somethingis =0?

OpenStudy (anonymous):

9x = -7y right?

OpenStudy (anonymous):

simplify it and plug it in either one of the equations

OpenStudy (anonymous):

and you can get one variable. then use ur simplified equation so get the other variable.

pooja195 (pooja195):

thats not right =/ 9x does not =7y this is confusing >_<

OpenStudy (anonymous):

9x is = to -7y because 7y must be the opposite of 9x to get 0

OpenStudy (anonymous):

get it?

pooja195 (pooja195):

No this is confusing now =/

pooja195 (pooja195):

Dont you have to have opposite symbols so that one can cancel out?

OpenStudy (anonymous):

that's the other choice but it's a lot harder

pooja195 (pooja195):

thats the way i was taught and thats how im going to be tested =/

OpenStudy (anonymous):

and since both are positive you cannot add

OpenStudy (anonymous):

you could only eliminate the other side, or the one that does not have a variable.

OpenStudy (anonymous):

so you wanna see the steps with explanation?

pooja195 (pooja195):

sure :)

hartnn (hartnn):

a+b = 0 then a = -b right ? so, if 9x+7y = 0 then 9x = -7y

pooja195 (pooja195):

I havent learned the way you guys are doing it =/

OpenStudy (anonymous):

4x+3y=-1 5x+4y=1 if you add them..... 9x+7y=0 and since it's 0 that must cancel out each other..... so you get 9x=-7y x=-7/9 y and so you plug it into one of the equations.... -7/9 y + 3y = -1 -4/9 y = -1 -4y = 9 -y = 9/4 y= -9/4 or y= -2.25 and you plug this back into the simplified equation... which is x=-7/9y x=-9/4 ( -7/9) x= 7/4 or x= 1.75 do you get it now? omg this is sooo long.....

pooja195 (pooja195):

O_O

OpenStudy (anonymous):

light?

OpenStudy (anonymous):

and the cancel out part is just moving y to the other side that's all.

pooja195 (pooja195):

I kind of get it =/

OpenStudy (anonymous):

oh yeah iggggnooooooooooooooorrrrrrreeeeeeeeeeeeee the answers!!!!! i had to type it in or else i can't solve it

OpenStudy (anonymous):

so which part do you NOT get?

pooja195 (pooja195):

ok i got it thanks so much :)

OpenStudy (anonymous):

ok..... phew. i hate those problems. u get messed up soooo easily

Nnesha (nnesha):

hmm very looong post well oaky pooja so you have to multiply one of equation to get one positive and one negative same number so yo cancel out that variable

OpenStudy (anonymous):

what? really? how old r u?

Nnesha (nnesha):

you can*

OpenStudy (anonymous):

oops am i interrupting you nnesha?

pooja195 (pooja195):

Thanks nnesha <3 i got it :)

OpenStudy (anonymous):

phew.

Nnesha (nnesha):

nope @Molly_Qian i was trying to explain the other method that she already knw :) thats all

OpenStudy (anonymous):

oh that one. it's a bit harder for me since there's more math. but ok.

Nnesha (nnesha):

yeah but the other one can be hard for her

OpenStudy (anonymous):

*him*

OpenStudy (anonymous):

well u can type it up and we can check answers